OpenStudy (anonymous):

What is the speed, in meters per second, of a paper airplane that flies 24 meters in 6 seconds? A. 144 B. 30 C. 18 D. 4

jimthompson5910 (jim_thompson5910):

speed = (distance)/(time) speed = (24 meters)/(6 seconds) speed = (24/6) m/s speed = 4 m/s

OpenStudy (anonymous):

The formula for speed is speed equals distance divided by time. S=D/T. So divide 24 by 6. so what is the answer

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes that is correct :)